OpenStudy (anonymous):

Simplify the expression 6C4

jimthompson5910 (jim_thompson5910):

use the formula n C r = (n!)/(r!*(n-r)!) in this case, you're plugging in n = 6 and r = 4

jimthompson5910 (jim_thompson5910):

Doing so will give you this n C r = (n!)/(r!(n-r)!) 6 C 4 = (6!)/(4!*(6-4)!) 6 C 4 = (6!)/(4!*2!) 6 C 4 = (6*5*4!)/(4!*2!) 6 C 4 = (6*5)/(2!) 6 C 4 = (6*5)/(2*1) I'll let you finish

OpenStudy (anonymous):

so the answer is 15. but I dont understand the excalmation point? and where you got 6 C 4 = (6*5*4!)/(4!*2!) the 5 and four next to the 6? and then you didnt multiply 4*2? i dont understand

OpenStudy (anonymous):

Its a combination. Hope you know that.

jimthompson5910 (jim_thompson5910):

something like 6! means 6*5*4*3*2*1

OpenStudy (anonymous):

! is a factorial symobol

jimthompson5910 (jim_thompson5910):

you multiply all the positive integers less than that value til you reach 1

jimthompson5910 (jim_thompson5910):

another example: 9! = 9*8*7*6*5*4*3*2*1

jimthompson5910 (jim_thompson5910):

the reason why I stopped at 4! is because there's a 4! already in the denominator so the two copies of 4! will cancel

jimthompson5910 (jim_thompson5910):

and you cannot say 2*4! = 8 this is because 4! isn't just 4, it's really 4*3*2*1 = 24

OpenStudy (anonymous):

ooooh!!!! i got it now! thank you !!

jimthompson5910 (jim_thompson5910):

you're welcome